An oxygen container that one can use in the mountains has a 90-min oxygen supply at a speed of

6 L/min (with the volume measured at atmospheric pressure).

A.Determine the volume of oxygen inside the container. Assume that the pressure inside the container is 6.9 × 106 N/m2 and the temperature is 20∘C.

Express your answer to two significant figures and include the appropriate units.

B.Determine the mass of the oxygen inside the container.

Express your answer to two significant figures and include the appropriate units.

Please show steps
Answer #1

A)

Here ,

time for oxygen , t = 90 min

volume flow rate = 6 L/min

volume of oxygen at atmospheric pressure , V = 6 * 90

V = 540 litres

Pressure , P2 = 6.9 *10^6 N/m^2

Using charles law ,

P1 * V1 = P2 * V2

540 * 1.01 *10^5 = 6.9 *10^6 * V2

V2 = 7.90 litres

the volume of oxygen container is 7.90 litres

B)

Now , using ideal gas equation

P*V = n*R*T

540 * 1 = n * 0.0821 * (20 + 273)

solving for n

n = 22.45 moles ,

mass of oxygen = 22.45 * 0.032

mass of oxygen = 0.7184 Kg

the mass of oxygen in the tank is 0.7184 kg